# Maël Dumas LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Anthony Perez LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France Mathis Rocton Algorithms and Complexity Group, TU Wien, Vienna, Austria Ioan Todinca LIFO EA 4022, Univ. Orléans, INSA Centre Val de Loire, France

### v2

@@ 1 Introduction
We focus here on a sub-family of parameterized algorithms, namely on kernelization. The goal of kernelization is to provide a polynomial algorithm transforming any instance $(I,k)$ of the problem into an equivalent instance $(I^{\prime},k^{\prime})$ where $k^{\prime}$ is upper bounded by some function of $k$ (in our case we will simply have $k^{\prime}\leqslant k$), and the size of the new instance $I^{\prime}$ is upper bounded by some function $g(k)$. Hence the size of the reduced instance does not depend on the size of the original instance. Kernels are obtained throught a set of reduction rules. While kernelization is possible for all FPT problems (the two notions are actually equivalent), the interesting question is whether a given FPT problem admits polynomial kernels, where the size of the reduced instance is bounded by some polynomial in $k$. Note that, under some complexity assumptions, not all FPT problems admit polynomial kernels [6, 7, 8, 11, 24, 31].
In this paper we focus on Block Graph Editing, Block Graph Deletion and all three variants of modification problems towards striclty chordal graphs. Block graphs are a well-studied subclass of chordal graphs, that are diamond-free (Figure 1). They are moreover the graphs in which every biconnected component induces a clique. Notice that this observation allows for a simple polynomial-time algorithm for Block Graph Completion since one needs to turn every biconnected component of the input graph into a clique. Strictly chordal graphs are another subclass of chordal graphs, also known as block duplicate graphs [28, 29, 23]. They can be obtained from block graphs by repeatedly choosing some cut-vertex $u$ and adding a true twin $v$ of $u$, that is a vertex $v$ adjacent to $u$ and all neighbors of $u$ [23]. They can also be characterized as dart, gem-free chordal graphs (see Figure 1 and next section) and are thus ptolemaic (i.e. chordal distance-hereditary [9]). Strictly chordal graphs are also known to be a subclass of 4-leaf power graphs [29], and a super-class of 3-leaf power graphs [16].
@@
Kernelization for chordal completion goes back to the ’90s and the seminal paper of Kaplan, Shamir and Tarjan [27]. In this work, the authors provide an $O(k^{3})$ vertex-kernel for Chordal Completion (also known as Minimum Fill-In) that was later improved to $O(k^{2})$ by a tighter analysis [35]. Since then, several authors addressed completion, deletion and/or editing problems towards sub-classes of chordal graphs, as 3-leaf power graphs [4], split and threshold graphs [25], proper interval graphs [5], trivially perfect graphs [3, 18, 21, 25, 19] or ptolemaic graphs [14]. All these classes have in common that they can be defined as chordal graphs, plus a constant number of obstructions. Several questions remain open, for example it is not known whether chordal deletion or chordal editing admit polynomial kernels [13]. If instead of modifying edges we want to remove vertices, the problem Block Graph Vertex Deletion has a kernel with $O(k^{4})$ vertices [1].
@@
We first prove that most variants of the considered problems are NP-Complete, the only exception being the Block Graph Completion problem that admits a polynomial-time algorithm. Secondly, we give kernelization algorithms for all these problems, with different bounds. First of all, we begin by illustrating the techniques we will use on the Block Graph Editing problem and obtain an $O(k^{2})$ vertex-kernel. This result can be transfered to Block Graph Deletion as well. Next, we present a kernelization algorithm for the Strictly Chordal Editing problem, producing a reduced instance with $O(k^{4})$ vertices. Eventually, we discuss how this approach leads to $O(k^{3})$ vertex-kernels for both Strictly Chordal Completion and Strictly Chordal Deletion. Above all, our purpose is to exhibit general techniques that might, we hope, be extended to kernelizations for edge modification problems towards other graph classes. Several such algorithms, e.g., [4, 21] share the following feature. Very informally, the target class $\mathcal{G}$ admits a tree-like decomposition, in the sense that the vertices of any graph $H\in{\mathcal{G}}$ can be partitioned into clique modules (having the same neighborhood in $H$), and these modules can be mapped onto the nodes of a decomposition tree, the structure of the tree describing the adjacencies between modules. Therefore, if an arbitrary graph $G$ can be transformed into graph $H$ by at most $k$ edge additions or deletions, at most $2k$ modules can be affected by the modifications. By removing the affected nodes from the decomposition tree, we are left with several components (chunks) that correspond, in the initial graph $G$ as well as in $H$, to induced subgraphs that may be large but that already belong to the target class. Moreover, these chunks are attached to the rest of graph $G$ in a very regular way, through one or two nodes of the decomposition tree. The kernelization algorithms need to analyze these chunks and provide reduction rules, typically by ensuring a small number of nodes in the decomposition tree, plus the fact that each node corresponds to a module of small size.
Both the classes of block graphs and of strictly chordal graphs do not have exactly a tree-like decomposition, but still can be decomposed into structures than can be seen as a generalization of a tree. Our algorithms exploit these informal observations and provide the necessary reduction rules together with the combinatorial analysis for the kernel size.
@@ 2 Preliminaries
Block graphs are graphs in which every biconnected component is a clique. They can also be characterized as chordal graphs that do not contain diamonds as induced subgraph [2] (see Figure 1). A natural generalization of block graphs are strictly chordal graphs, also known as block duplicate graphs, that are obtained from block graphs by adding true twins [28, 29, 23].
Theorem 1 ([33]).
@@
    The critical clique graph $\mathcal{C}(G)$ is a block graph,
@@
Let $\mathcal{G}$ be an hereditary class of graphs closed under true twin addition. For every graph $G=(V,E)$, there exists an optimal edition (resp. completion, deletion) $F$ into a graph of $\mathcal{G}$ such that for any two critical cliques $K$ and $K$’ either $(K\times K^{\prime})\subseteq F$ or $(K\times K^{\prime})\cap F=\emptyset$.
@@
From Lemma 2, $G$ is chordal and its minimal separators of are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ or $S_{2}$ if they are not maximal cliques. If $S_{i}$ intersects exactly one maximal clique, $i\in\{1,2\}$, it is clear that it does not intersect any minimal separator of $G_{i}$. If $S_{i}$ is a critical clique, we claim that either $S_{i}$ is a minimal separator of $G_{i}$ or does not intersect any minimal separator of $G_{i}$. Indeed, if there is a minimal (clique) separator $S$ of $G_{i}$ such that $S_{i}\subset S$, then since $S_{i}$ is a critical clique, there exist $u\in S_{i},v\in S\setminus S_{i}$ and $w\in V(G_{i})\setminus S$ such that $w$ is adjacent to exactly one of $u$ and $v$. Suppose w.l.o.g. that $w$ is adjacent to $u$ and not to $v$, then $u$ is in a minimal $vw$-separator that intersects $S$ and is not equal to it. This is a contradiction since $G_{i}$ is strictly chordal and from Theorem 1 its minimal separators are pairwise disjoint. It follows that the minimal separators of $G$ are pairwise disjoint and by Theorem 1 $G$ is strictly chordal. ∎
@@
The following result is crucial to bound the size of the kernels for all variants of the studied modification problems.
@@
Let $G=(V,E)$ be a connected block graph, a set $A\subseteq V(G)$ and $T_{A}$ be a minimal connected induced subgraph of $G$ that spans all vertices of $A$. Denote by $f(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$. The following holds:
@@
    The graph $T_{A}\setminus(A\cup f(T_{A}))$ contains at most $2\cdot|A|$ connected components.
@@
A convenient way to represent the tree-like structure of block graph $G$ is its block-cut tree $T_{BC}(G)$. Recall that the block-cut tree of a graph $G$ has two types of nodes: the block nodes correspond to blocks of $G$ (i.e, biconnected components which, in our case, are precisely the maximal cliques of $G$) and the cut nodes correspond to cut-vertices of $G$. We put an edge between a cut node and a block node in $T_{BC}(G)$ if the corresponding cut-vertex belongs to the corresponding block of $G$.
Each vertex of $G$ that is not a cut-vertex belongs to a unique block of $G$. Therefore we can map each vertex $v$ of $V(G)$ on a unique node $n(v)$ of $T_{BC}(G)$ as follows: if $v$ is a cut-vertex, we map it on its corresponding cut node in $T_{BC}(G)$, otherwise we map it on the block node corresponding to the unique block of $G$ containing $v$. Observe that, for any two vertices $u,v\in A$, any (elementary) path $P$ from $u$ to $v$ in $G$ corresponds to the path $P_{BC}$ from $n(u)$ to $n(v)$ in $T_{BC}(G)$. In particular, $P$ contains all vertices corresponding to cut nodes of $P_{BC}$, therefore $T_{A}$ must contain all these vertices. Altogether, the vertices of $T_{A}$ are precisely the vertices of $A$, plus all cut-vertices of $G$ corresponding to the cut nodes of such paths, implying the unicity of $T_{A}$.
Let $T_{BC}(A)$ denote the subtree of $T_{BC}(G)$ spanning all nodes of $n(A)=\{n(a)\mid a\in A\}$. We count the vertices of $f(T_{A})\setminus A$. If $|A|=2$ then $f(T_{A})\setminus A=\emptyset$. Suppose now that $|A|\geqslant 3$ and let $a^{\prime}\in f(T_{A})\setminus A$. By the previous observation, it is a cut-vertex of $G$, so $n(a^{\prime})$ is a cut node of $T_{BC}(A)$. Let $v$ be a neighbor of $a^{\prime}$ in $T_{A}$. By construction of $T_{A}$ and $T_{BC}(A)$, there is a block node $b$ adjacent to $n(a^{\prime})$ in $T_{BC}(A)$, such that $v$ and $a^{\prime}$ are in the maximal clique of $G$ corresponding to node $b$. Moreover, $v$ is in $A$, or $v$ is a cut-vertex of $G$ such that the cut node $n(v)$ is adjacent to $b$ in $T_{BC}(A)$. Hence we have:
  1. 1.
    $n(a^{\prime})$ is of degree at least 3 in $T_{BC}(A)$, or
  2. 2.
    $n(a^{\prime})$ is the neighbor of a block node $b$ of degree at least three in $T_{BC}(A)$, or, if none of these hold, then
  3. 3.
    $n(a^{\prime})$ has exactly two neighbors $b$ and $b^{\prime}$ in $T_{BC}(A)$, the corresponding maximal cliques of $G$ contain at least $p$ vertices of $A$, where $p$ plus the degrees of $b$ and $b^{\prime}$ in $T_{BC}(A)$ is at least three.
Let $k$ be the number of leaves of $T_{BC}(A)$. Observe that for any leaf $l$ of $T_{BC}(A)$, there is some $a\in A$ such that $n(a)=l$. Choose for each leaf $l$ a unique vertex $a_{l}\in A$ such that $n(a_{l})=l$, we call $a_{l}$ a leaf vertex. Note that the number of vertices $a^{\prime}\in f(T_{A})$ corresponding to the first two items of the above enumeration is upper bounded by $3k$. Indeed, $n(a^{\prime})$ is incident to an edge of $T_{BC}(A)$, having an end node of degree at least 3. One can easily check that, in any tree of $k$ leaves, the number of such edges is bounded by $3k$ (this can be shown by induction on the number of leaves of the tree, adding a new leaf node at a time). We count now the vertices $a^{\prime}\in f(T_{A})$ of the third type. By the third item, $a^{\prime}$ has at least one neighbor $a\in A$ in graph $T_{A}$, such that $a$ is not a leaf vertex. Observe that $a$ can be in the neighborhood of at most two vertices $a^{\prime}\in f(T_{A})$ of this third type. Altogether it follows that $|f(T_{A})\setminus A|\leqslant 3k+2(|A|-k)\leqslant 3\cdot|A|$.
To prove the third item, the number of components of $T_{A}\setminus(A\cup f(T_{A}))$, we visualize again the situation in $T_{BC}(A)$. Recall that any node of degree at least three in $T_{BC}(A)$ is either a cut node, in which case it is in $f(T_{A})$ (the first case of the enumeration above) or a block node, but then all its neighbors in $T_{BC}(A)$ correspond to vertices of $f(T_{A})$ (the second case of the enumeration above). Therefore the components of $T_{A}\setminus f(T_{A})$ correspond to the components of $T_{BC}(A)$ after removal of all nodes of degree at least $3$. Hence the number of such components is upper bounded by $2k$. Removing the leaf vertices of $A$ does not increase the number of components, and the removal of each other vertex of $A$ increases the number of components by at most one. Thus $T_{A}\setminus(A\cup f(T_{A}))$ has at most $2\cdot|A|$ components, concluding the proof of our lemma. ∎

### v3

@@ 1 Introduction
We focus here on a sub-family of parameterized algorithms, namely on kernelization. The goal of kernelization is to provide a polynomial algorithm transforming any instance $(I,k)$ of the problem into an equivalent instance $(I^{\prime},k^{\prime})$ where $k^{\prime}$ is upper bounded by some function of $k$ (in our case we will simply have $k^{\prime}\leqslant k$), and the size of the new instance $I^{\prime}$ is upper bounded by some function $g(k)$. Hence the size of the reduced instance does not depend on the size of the original instance. Kernels are obtained through a set of reduction rules. While kernelization is possible for all FPT problems (the two notions are actually equivalent), the interesting question is whether a given FPT problem admits polynomial kernels, where the size of the reduced instance is bounded by some polynomial in $k$. Note that, under some complexity assumptions, not all FPT problems admit polynomial kernels [5, 6, 7, 10, 25, 33].
In this paper we focus on Block Graph Editing, Block Graph Deletion and all three variants of modification problems towards strictly chordal graphs. Block graphs are a well-studied subclass of chordal graphs, that are diamond-free (Figure 2). They are moreover the graphs in which every biconnected component induces a clique. Notice that this observation allows for a simple polynomial-time algorithm for Block Graph Completion since one needs to turn every biconnected component of the input graph into a clique. Strictly chordal graphs are another subclass of chordal graphs, also known as block duplicate graphs [30, 31, 24]. They can be obtained from block graphs by repeatedly choosing some cut-vertex $u$ and adding a true twin $v$ of $u$, that is a vertex $v$ adjacent to $u$ and all neighbors of $u$ [24]. They can also be characterized as dart, gem-free chordal graphs (see Figure 2 and next section) and are thus ptolemaic (i.e. chordal distance-hereditary [8]). Strictly chordal graphs are also known to be a subclass of 4-leaf power graphs [31], and a super-class of 3-leaf power graphs [16].
@@
Kernelization for chordal completion goes back to the ’90s and the seminal paper of Kaplan, Shamir and Tarjan [29]. In this work, the authors provide an $O(k^{3})$ vertex-kernel for Chordal Completion (also known as Minimum Fill-In) that was later improved to $O(k^{2})$ by a tighter analysis [37]. Since then, several authors addressed completion, deletion and/or editing problems towards sub-classes of chordal graphs, such as 3-leaf power graphs [3], split and threshold graphs [26], proper interval graphs [4], trivially perfect graphs [2, 19, 22, 26, 20] or ptolemaic graphs [13]. We provide a general picture for some of those classes Figure 1.
chordal [29, 37]$4$-leaf powerintervalptolemaic [13]trivially perfect [2, 20]strictly chordalsplit [2, 27]threshold [18]proper interval [4]$3$-leaf power [3]blockpoly-kernel $\downarrow$ CompletionEditing, Deletion
Figure 1: Kernelization status of subclasses of chordal graphs. A class below the red or blue line indicates that the corresponding completion, deletion or editing problem admits a polynomial kernel. All the edge modification problems for the presented classes are NP-complete at the exception of the ones for $4$-leaf power for which the complexity is unknown, Split Editing which is surprisingly in P [27], and Block Graph Completion which is also in $P$ as observed previously.
All these classes have in common that they can be defined as chordal graphs, plus a constant number of obstructions. Moreover, several known results deal with classes that share a very common feature (see e.g. [4, 32, 3, 20]). Very informally, the target class $\mathcal{G}$ admits a tree-like decomposition, in the sense that the vertices of any graph $H\in{\mathcal{G}}$ can be partitioned into clique modules (having the same neighborhood in $H$), and these modules can be mapped onto the nodes of a decomposition tree, the structure of the tree describing the adjacencies between modules. Therefore, if an arbitrary graph $G$ can be transformed into graph $H$ by at most $k$ edge additions or deletions, at most $2k$ modules can be affected by the modifications. By removing the affected nodes from the decomposition tree, we are left with several components (chunks) that correspond, in the initial graph $G$ as well as in $H$, to induced subgraphs that may be large but that already belong to the target class. Moreover, these chunks are attached to the rest of graph $G$ in a very regular way, through one or two nodes of the decomposition tree. The kernelization algorithms need to analyze these chunks and provide reduction rules, typically by ensuring a small number of nodes in the decomposition tree, plus the fact that each node corresponds to a module of small size. However, finding a general framework that captures this particular decomposition does not seem to be an easy task as each class exhibits particular structural properties that may need to be dealt with independently.
As mentioned earlier, when restricted to completion problems any input graph cannot have large induced cycles and this property may be exploited to design kernelization algorithms. This observation lead to the following conjecture by Bessy and Perez [4].
Conjecture 1 ([4]).
Let $\mathcal{G}$ be a class of graphs that can be defined as chordal graphs plus a constant number of obstructions Then, the $\mathcal{G}$-Completion problem admits a polynomial kernel.
Several other questions remain open, for example it is not known whether chordal deletion or chordal editing admit polynomial kernels [12].
@@
We first prove that most variants of the considered problems are NP-Complete, the only exception being the Block Graph Completion problem that admits a polynomial-time algorithm. Secondly, we give kernelization algorithms for all these problems, with different bounds. First of all, we begin by illustrating the techniques we will use on the Block Graph Editing problem and obtain the following result.
Theorem 1.
Block Graph Editing and Block Graph Deletion admit a kernel with $O(k^{2})$ vertices.
Next, we present kernelization algorithms for the Strictly Chordal Editing problem and obtain better vertex-kernels for both Strictly Chordal Completion and Strictly Chordal Deletion.
Theorem 2.
Strictly Chordal Editing admits a kernel with $O(k^{4})$ vertices.
Theorem 3.
Above all, our purpose is to exhibit general techniques that might, we hope, be extended to kernelizations for edge modification problems towards other graph classes. In particular, unlike the general strategy we previously described, both the classes of block graphs and of strictly chordal graphs do not have exactly a tree-like decomposition. Still, they can be decomposed into structures than can be seen as a generalization of a tree. Our algorithms exploit these informal observations and provide the necessary reduction rules together with the combinatorial analysis for the kernel size. Finally, we note that while both kernels use structures and rules that are similar, the ones for Block Graph Editing are significantly simpler, hence the kernel for this problem is a nice warm-up for the one for Strictly Chordal Editing. Moreover, even though reduction rules for Strictly Chordal Editing may be directly adapted for Block Graph Editing, it would lead to a significantly larger kernel.
@@ 2 Preliminaries
Block graphs are graphs in which every biconnected component is a clique. They can also be characterized as chordal graphs that do not contain diamonds as induced subgraph [1] (see Figure 2). A natural generalization of block graphs are strictly chordal graphs, also known as block duplicate graphs, that are obtained from block graphs by adding true twins [30, 31, 24]. Note that by definition the critical clique graph $\mathcal{C}(G)$ is a block graph.
Theorem 4 (Theorem 8 [35]).
@@
Let $\mathcal{G}$ be an hereditary class of graphs closed under true twin addition. For every graph $G=(V,E)$, there exists an optimal edition (resp. completion, deletion) $F$ into a graph of $\mathcal{G}$ such that for any two critical cliques $K$ and $K^{\prime}$ either $(K\times K^{\prime})\subseteq F$ or $(K\times K^{\prime})\cap F=\emptyset$.
@@
From Lemma 2, $G$ is chordal and its minimal separators are the ones of $G_{1}$ and $G_{2}$ plus possibly $S_{1}$ or $S_{2}$ if they are not maximal cliques. If $S_{i}$ intersects exactly one maximal clique, $i\in\{1,2\}$, it is clear that it does not intersect any minimal separator of $G_{i}$. If $S_{i}$ is a critical clique, we claim that either $S_{i}$ is a minimal separator of $G_{i}$ or does not intersect any minimal separator of $G_{i}$. Indeed, if there is a minimal (clique) separator $S$ of $G_{i}$ such that $S_{i}\subset S$, then since $S_{i}$ is a critical clique, there exist $u\in S_{i},v\in S\setminus S_{i}$ and $w\in V(G_{i})\setminus S$ such that $w$ is adjacent to exactly one of $u$ and $v$. Suppose w.l.o.g. that $w$ is adjacent to $u$ and not to $v$, then $u$ is in a minimal $vw$-separator that intersects $S$ and is not equal to it. This is a contradiction since $G_{i}$ is strictly chordal and from Theorem 4 its minimal separators are pairwise disjoint. It follows that the minimal separators of $G$ are pairwise disjoint and by Theorem 4 $G$ is strictly chordal. ∎
@@
The following result is crucial to bound the size of the kernels for all variants of the studied modification problems. For a graph $G$, let $d(G)$ be the set of vertices of degree at least $3$ in $G$.
@@
Let $G=(V,E)$ be a connected block graph, $A$ be a subset of $V(G)$ and $T_{A}$ be a minimal connected induced subgraph of $G$ that spans all vertices of $A$. Denote by $d(T_{A})$ the set of vertices of degree at least $3$ in $T_{A}$. The following holds:
@@
    The graph $T_{A}\setminus(A\cup d(T_{A}))$ contains at most $2\cdot|A|$ connected components.
@@
We first prove the uniqueness of $T_{A}$. Observe that the internal vertices of a shortest path $P$ in $G$ between two vertices $u$ and $v$ are cut-vertices of $G$, since it is a block graph. Moreover, for any such internal vertex $w$ of $P$, vertices $u$ and $v$ are in different components of $G-w$. Therefore, $P$ is unique, and any path between $u$ and $v$ must contain every vertex of $P$. Hence, the vertices of $T_{A}$ are either in $A$ or separating vertices of $A$, and the uniqueness of $T_{A}$ follows, proving Item i.
To prove Item ii we will proceed by induction on the size of $A$. For $|A|\leqslant 2$ we have that $|d(T_{A})|=0$. Suppose now that $|A|\geqslant 3$. Let $a\in A$ be a non-cut-vertex of $T_{A}$. Such a vertex exists, e.g., it is sufficient to choose a longest path of $T_{A}$, and let $a$ be one of the endpoints. Then $a$ is not a cut vertex of $T_{A}$ (as the end of a longest path), and $a\in A$ by minimality of $T_{A}$. Notice moreover that $a$ is in only one maximal clique of $T_{A}$. Let $A^{\prime}=A\setminus\{a\}$ and let $T_{A^{\prime}}$ be a minimal connected induced subgraph of $G$ that spans all vertices of $A^{\prime}$. Observe that $T_{A^{\prime}}$ is an induced subgraph of $T_{A}$. We show that $|d(T_{A})|\leqslant|d(T_{A^{\prime}})|+3$. Since $T_{A}$ is connected, there exists a path in $T_{A}$ from $a$ to a vertex $u\in V(T_{A})\setminus V(T_{A^{\prime}})$ adjacent to some vertices of $T_{A^{\prime}}$ ($a=u$ is possible), moreover $u$ is unique. Any vertex on the path except for $u$ is of degree at most $2$ in $T_{A}$. Let $N=N_{G}(u)\cap V(T_{A^{\prime}})$ and note that since $T_{A^{\prime}}$ is a connected block graph, $N$ is a clique. By construction, the vertices of $T_{A^{\prime}}$ have the same degree in $T_{A}$ and $T_{A^{\prime}}$, except for the ones in $N$. Since $u$ might be of degree $3$ or more, if $|N|\leqslant 2$ then $|d(T_{A})|\leqslant|d(T_{A^{\prime}})|+3$. If $|N|\geqslant 4$, then all vertices in $N$ already have degree $3$ in $T_{A^{\prime}}$, hence $|d(T_{A})|=|d(T_{A^{\prime}})|+1$. If $|N|=3$ and $N$ contains a vertex of degree $3$ in $T_{A^{\prime}}$ then $|d(T_{A})|\leqslant|d(T_{A^{\prime}})|+3$. Otherwise, we claim that $V(T_{A^{\prime}})=N$. Indeed, if all vertices in $N$ have degree $2$ in $T_{A^{\prime}}$, the only possibility is that $N$ is exactly the vertex set of $T_{A^{\prime}}$, because $T_{A^{\prime}}$ is connected, $N$ induces a clique of size $3$ in this graph and no vertex of $N$ has other neighbors than the ones in $N$. In this case we have $|A|=4$ and $|d(T_{A})|=4$, which satisfies Item ii. Therefore, except for the latter case, we have $|d(T_{A})|\leqslant|d(T_{A^{\prime}})|+3$, and using the induction hypothesis we have $|d(T_{A})|\leqslant|d(T_{A^{\prime}})|+3\leqslant 3\cdot|A^{\prime}|+3=3\cdot|A|$.
We are left with Item iii, which can be proven in a similar way as the previous one. Using the same notation, the only vertices of $T_{A^{\prime}}$ whose degree may increase in $T_{A}$ are in $N$. Therefore, every connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ that does not contain vertices of $N$ is also a connected component of $T_{A}\setminus(A\cup d(T_{A}))$. Moreover, a connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ may contain vertices of $N$ only if these vertices have degree at most $2$ in $T_{A^{\prime}}$ and are not in $A$. In particular $|N|\leqslant 2$and both vertices lie in the same connected component of $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$ (if any). Therefore, this connected component may be split in two disjoint ones in $T_{A}\setminus(A\cup d(T_{A}))$ if the vertices of $N$ are removed. Moreover, an additional connected component corresponding to the path between $a$ and $u$ may exist in $T_{A}$. Hence, we can only increase the number of connected components of $T_{A}\setminus(A\cup d(T_{A}))$ by compared to $T_{A^{\prime}}\setminus(A^{\prime}\cup d(T_{A^{\prime}}))$. Since the property is verified for $|A|=1$, it follows by induction that there are $2\cdot|A|$ connected components in $T_{A}\setminus(A\cup d(T_{A}))$. ∎